### v4

# thanks: Present address

Abstract
This study investigates the phenomenon of the early-stage impact dynamics in dense suspensions. While Stokes flow theory traditionally applies to suspensions with micrometer-sized particles due to their low Reynolds numbers, suspensions containing larger particles defy such idealizations. Our work focuses on the early-stage impact dynamics in suspensions containing millimeter-sized particles through experiments of impact dynamics. We are particularly interested in the maximum drag force $F_{\mathrm{max}}$ acting on the projectile as a function of the impact speed $u_{0}$. We have successfully conducted experiments using these suspensions and confirmed the relation $F_{\mathrm{max}}\sim u_{0}^{3/2}$ for relatively large $u_{0}$ as observed in the previous studies of suspensions of micrometer-sized particles. Our findings reveal that the early-stage behaviors of millimeter-sized particle suspensions align well with predictions from the floating model, which is equivalent to the viscous model, typically applicable under Stokes flow conditions. This research sheds light on the impact dynamics in dense suspensions, particularly with larger particles, advancing our understanding beyond conventional micrometer-sized systems.
preprint: AIP/123-QED
I Introduction
The investigation of impact dynamics within dense suspensions has garnered substantial interest due to its broad applicability across industrial processes and planetary sciences [1]. An intriguing phenomenon within dense suspensions is impact-induced hardening (IIH), enabling locomotion to stop liquids while causing sinking for individuals standing or walking [2]. This behavior finds practical utility in applications such as protective vests [3]. Extensive experimental [4, 5, 6, 7, 8, 9] and theoretical [10, 11, 12] studies have explored IIH, distinguishing it from discontinuous shear thickening (DST) [2, 13, 14, 15] by its localized nature versus DST’s global manifestation, transient behavior versus steadiness, and distinct shear stress responses [10].
A primary focus in impact dynamics research involves elucidating the impact experiments of projectiles[16, 17, 18], where the time evolution of projectiles can be observed to determine the force based on contact mechanics[19]. Early-stage impact dynamics in dense suspensions is governed initially by a viscous force [9, 11, 12], and subsequently, an elastic force plays an important role [8, 10, 11, 12]. Notably, empirical observations suggest a relationship between the maximum drag force $F_{\mathrm{max}}$ and impact speed $u_{0}$, approximated by $F_{\mathrm{max}}\sim u_{0}^{3/2}$ predicted by the viscous model [9] and the floating model [11, 12] which are equivalent models. This result is similar to the relation $F_{\mathrm{max}}\sim u_{0}^{4/3}$ observed for impact processes in dry granular materials [20, 21].
The previous theoretical studies [10, 11, 12] rely on a lattice-Boltzmann method (LBM) and a discrete element method. Although their model analysis agrees with that of experimental results [9], there are various questions about the applicability of their analysis. Their approach is believed to be valid only if the suspended particles are immersed in a viscous fluid in which the Reynolds number is quite low and the Stokes approximation is valid. Because of the limitation of their computer resources, their simulation contains only less than 3000 suspension particles in fluids, and the radius of the projectile is only a few times larger than that of the suspension particles. If suspended particles are a typical suspension, such as cornstarch or potato starch, the projectile must be quite small. Increasing the size of the projectile while maintaining a constant size ratio between the suspension particles and the projectile may invalidate the Stokes approximation at the millimeter scale. This observation highlights the importance of the size scale of the dispersed particles in understanding impact processes at this scale.
Building upon prior investigation, this study experimentally demonstrates how the size scale of dispersed particles influences the early-stage dynamics of the impact processes in dense suspensions. Our exploration delves deeper into the mechanical response of suspension materials to sudden mechanical stimuli, uncovering the fundamental physical mechanisms. We critically reassess the applicability of theoretical analyses utilizing the floating model and LBM within our experimental framework, mainly focusing on early-stage dynamics and the relationship between $F_{\mathrm{max}}$ and $u_{0}$. Our findings affirm that the floating model or the viscous model, typically valid under Stokes flow conditions, holds relevance even for dense suspensions containing millimeter-sized particles.
The rest of the paper is structured as follows. In the next section, we depict the experimental setup. We introduce the theoretical framework, the floating model in detail in Sec. III. We explain the experimental results to clarify the applicability of the floating model in Sec. IV. In Sec. V we discuss our results. We conclude our results with some remarks in Sec. VI. In Appendix A, we clarify the sidewall effect used in the main text from the comparison of the experimental result with a cube container. In Appendix B, we examine whether the simplified buoyancy force can be used to describe experimental results. In Appendix C, we present the exact solution of the floating model.
II Experiment
Figure 1: An illustration of our experimental setup. We prepare a suspension containing millimeter-sized bi-disperse particles (6.0, 8.0 mm) in a solvent of NaCl solution confined in a quasi-two-dimensional container. The density of the solvent is matched with that of the particles. By dropping a metallic spherical projectile from the heights $h$, the projectile collides with the suspension liquid with the impact speed $u_{0}$. We record the impact processes in a high-speed camera.
Let us explain the setup of our experiment. We prepare a dense suspension by mixing millimeter-sized bi-disperse particles, consisting of two types of airsoft pellets: one with a maximum diameter of $d_{\rm max}=8.0$ mm and a density of $1.03~{}{\rm g/cm^{3}}$ (MARUSHIN KOGYO CO.), and the other with a minimum diameter of $d_{\rm min}=6.0$ mm and a density of $1.04~{}{\rm g/cm^{3}}$ (Tokyo Marui Co.), in a NaCl solution whose density is nearly matched with that of the dispersed particles ($\rho_{\rm f}\approx 1.05~{}{\rm g/cm^{3}}$). To observe the impact processes, we primarily use suspensions confined in a quasi-two-dimensional container ( Kenis, Ltd.) with the width $W=123{\rm mm}$, height $H=112{\rm mm}$, and depth $D=14{\rm mm}$.   [^1] As shown in Appendix A, we also examine a cube system with $W=D=94$mm, $H=90$mm to study the boundary effects in the quasi-two-dimensional container. See Fig. 1 as an illustration of the experimental setup.
We introduce the designed packing fraction defined as $\varphi:=V_{\rm p}/V_{\rm s}$ to characterize the density of $N$ number of each particle, where $V_{\rm p}:=\frac{4}{3}\pi N\left[\left(d_{\rm max}/2\right)^{2}+\left(d_{\rm min}/2\right)^{2}\right]$ and $V_{\rm s}:=W\times D\times H$. We prepare the suspension liquid in four different designed packing fractions $\varphi=0.40,0.48,0.51,$ and $0.56$, which correspond to the number of particles $N=243,291,310,$ and $338$ respectively. Note that particles are not uniformly distributed, and most of them float to form a densely packed cluster due to the incomplete density matching. We use the effective packing fraction $\phi$ in the floating cluster near the surface of water for our data analysis in Sec. IV, where $\varphi=0.40,0.48,0.51$, and $0.56$ correspond to $\phi=0.52,0.53,0.54$, and $0.56$, respectively.
We use a metallic sphere as the projectile (diameter $D_{\rm I}=12$ mm, $\rho_{\rm I}=7800~{}{\rm kg\cdot m^{-3}}$), which is suspended by an electromagnet (ESCO Co., Ltd., EA984CM-1) above the suspension before impact. After switching off the magnetic force, the projectile is dropped into the suspension with an initial impact speed $u_{0}$. The impact process is recorded using a high-speed camera (Phantom V641, Phantom T1340, Vision Research) at a frame rate of 10,000 fps. The impact velocity is controlled by varying the release height of the projectile ($h=4\sim 720~{}{\rm mm}$). We conduct repeated impact experiments (10 to 16 times per height) to minimize specific responses from peculiar configurations of the suspended particles.
We extract the trajectory of the projectile by an optical tracking technique with the aid of the Open CV library of Python. We set the marker on the top of the projectile to get its position in each flame. By differentiating the time-evolution of the position with respect to time, we evaluate the time-evolutions of the velocity and the acceleration of the projectile. Subtracting the buoyancy force acting on the projectile from the equation of motion of the projectile, we can evaluate the drag force acting on the projectile.
III Floating model
Brassard et al. [9] proposed the viscous model to describe the motion of the projectile in dense suspensions of micron-meter-sized particles in the early stage of the impact dynamics. Later, Pradipto and Hayakawa [11, 12] renamed it as the floating model, adding its derivation. To get IIH, we need an elastic force [6, 7, 8, 10, 11, 12], which appears only in a relatively late stage of the impact process, although the floating model involves only the viscous drag force acting on the projectile, except for the buoyancy force. In this paper, we adopt the floating model to describe the motion around $t_{\mathrm{max}}$ at which the drag force takes $F_{\mathrm{max}}$. We depict an impact process in a suspension in the left figure of Fig. 2, where the position of the bottom head of the projectile with radius $a_{\rm I}$ and its density $\rho_{\rm I}$ is denoted by $z$ and $z=0$ is fixed on the surface of the suspension. Thus, we are only interested in the position $z$ satisfying $z\leq 0$, i,e, under the influence of the suspension liquid. As a result of the drag force $F_{\rm D}$ and buoyancy force $F_{\mathrm{B}}$ acting on the projectile, the projectile speed decreases with time and approaches zero. We should note that the maximum force $F_{\mathrm{max}}$ acting on the projectile appears in the relatively early stage of impact processes, and the behavior of the projectile around $t_{\mathrm{max}}$ can be understood without the elastic force acting on the projectile, at least, for micro-meter sized suspensions and LBM simulations [9, 11, 12].
Figure 2: An illustration of a projectile in a suspension liquid, where the impact speed is $u_{0}$ when the projectile attaches to the surface of the liquid (Left). We denote the position of the bottom head of the projectile $z(<0)$ where the center of mass is located at $z_{\rm I}$, the radius of the projectile $a_{\rm I}$, and the density of it is $\rho_{\rm I}$. The density of the solvent is $\rho_{\rm f}$. We introduce the effective viscosity $\eta_{\mathrm{eff}}$ acting on the projectile as a mean-field viscosity of the suspension liquid in the floating model (Right).
Let us discuss the buoyancy force $F_{\mathrm{B}}$ acting on the projectile. Although $F_{\mathrm{B}}$ should be proportional to the volume inside the liquid and the gravity directly acts on the projectile in the part outside the liquid, we simply adopt an approximate expression for $F_{\mathrm{B}}$ as [11, 12]
$F_{\rm B}\simeq-m_{\rm I}g\left(1-\frac{\rho_{\rm f}}{\rho_{\rm I}}\right).$ (1)
where $m_{\rm I}$ is the mass of the projectile, $g$ is the gravitational constant, $\rho_{\rm f}$ is the density of the solvent and $\rho_{\rm I}$ is the density of the projectile. The validity of this simplification is argued in Appendix B.
Next, we discuss the drag force $F_{\mathrm{D}}$ acting on the projectile. In the floating model [11, 12] as well as the viscous model [9], we assume that the drag force is proportional to the moving velocity. This drag force is the result of a mean-field approximation of the suspension liquid as depicted in the right figure of Fig. 2. Note that the derivation of the floating model can be found in Ref. [11]. If we assume such a model, the drag force acting on a partially filled projectile in the liquid is expressed as
$F_{\rm D}=3\pi\eta_{{\rm eff}}\frac{dz}{dt}z$ (2)
where $\eta_{\rm eff}$ is an effective viscosity. For simplicity, we assume that $\eta_{\mathrm{eff}}$ is independent of $u_{0}$ but depends on $\phi$. The expression of Eq. (2) is reduced to the Stokes drag force for a filled projectile in the liquid for $z\leq-2a_{\rm I}$. Thus, Eq. (2) is only valid for $-2a_{\rm I}<z<0$. It is controversial whether the drag force $F_{\mathrm{D}}$ can be used for a suspension of millimeter-sized particles because fluid flows around a large particle to generate vortices. Nevertheless, we may examine Eq. (2) for a fluid flow around a projectile in a dense suspension of millimeter-sized particles because the interstitial distance between suspended particles is too small to generate vortices. The validity of this model will be examined by the comparison between the model analysis and experimental results.
With the aid of Eqs. (1) and (2), the motion of the projectile in the early stage impact dynamics may be described by
$m_{\rm I}\frac{d^{2}z}{dt^{2}}=-m_{\rm I}g\left(1-\frac{\rho_{\rm f}}{\rho_{\rm I}}\right)+3\pi\eta_{{\rm eff}}\frac{dz}{dt}z.$ (3)
The dimensionless form of Eq. (3) is expressed as
$\frac{d^{2}Z}{d\tau^{2}}=-\left(1-\xi\right)+\eta\frac{dZ}{d\tau}Z$ (4)
where
$Z:=\frac{z}{a_{\rm I}},~{}\tau:=\sqrt{\frac{g}{a_{\rm I}}}t,~{}\xi:=\frac{\rho_{\rm f}}{\rho_{\rm I}},~{}\eta:=\frac{3\pi\eta_{{\rm eff}}a_{\rm I}^{3/2}}{g^{1/2}m_{\rm I}},~{}U_{0}:=\frac{u_{0}}{\sqrt{a_{\rm I}g}}.$ (5)
Equation (4), which is equivalent to Eq. (3), is the equation of the floating model [9, 11, 12].
As shown in Refs. [11, 12] and Appendix C, we can solve Eq. (4) exactly. The exact solution of Eq. (4) leads to the power-law relations of $F_{\mathrm{max}}$ with the high impact speed $U_{0}$[11]
$\tilde{F}_{\mathrm{max}}=\frac{\sqrt{2\eta}}{3}U_{0}^{\frac{3}{2}},\qquad\left(\eta U_{0}\gg 1\right)$ (6)
where $\tilde{F}_{\mathrm{max}}:=F_{\mathrm{max}}/m_{\rm I}g$.
IV Results
Figure 3: A set of experimental snapshots of an impact process for $\phi=0.56$ and $u_{0}=3.1~{}{\rm m/s}$. Note that 2.1 ms after the first impact (the middle figure) the force acting on the projectile exhibits the maximum value $F_{\mathrm{max}}$. The movie is available at Ref.[22].
Let us present the experimental results and examine whether the floating model introduced in the previous section can be used in the early stage of dynamics focusing on the relationship between $F_{\mathrm{max}}$ and $u_{0}$. Figure 3 is a typical set of experimental snapshots of an impact process for $\varphi=\phi=0.56$ and $u_{0}=3.1~{}{\rm m/s}$. The impact speed $u_{0}$ and the initiation time of the impact ($t=0$) are defined as the points of its maximum velocity and the corresponding time, respectively. Then, the time $t$ begins with the initiation time. Figure 3 (b) is a snapshot corresponding to the first hit of the projectile on the liquid surface at $t=0$. Figure  3 (c) is a snapshot when the drag force acting on the projectile reaches the maximum value $F_{\mathrm{max}}$ at $t_{\mathrm{max}}=2.1$ ms after the first impact. The movie of the impact process is available in Ref. [22].
Figure 4: The time evolution of the velocity of the projectile (a) and the drag force acting on the projectile (b) for $\phi=0.56$ and $u_{0}=3.1~{}{\rm m/s}$. The blue dots are experimental data. The solid lines are the solution of Eq. (3) with a fitting parameter $\eta$. The dashed vertical red line indicates the time $t_{\mathrm{max}}$ to take $F_{\mathrm{max}}$.
The projectile stops in the middle of the suspension for the dense suspensions ($\phi\geq 0.54$ corresponding to $\varphi\geq 0.51$), where the projectile cannot sneak in the suspension liquid for low-speed impact and it can sneak in the liquid completely for high-speed impact. For lower packing fractions such as $\phi=0.52$ (corresponding to $\varphi=0.40$), the projectile penetrates the liquid completely to reach the bottom of the container.
Figure 5: The plots between the dimensionless maximum force $\tilde{F}_{\mathrm{max}}:=F_{\mathrm{max}}/m_{\rm I}g$ and impact-speed $U_{0}:=u_{0}/\sqrt{a_{\rm I}g}$ for various packing fractions $\phi=0.52$ (a), 0.53 (b), 0.54 (c), and 0.56 (d) corresponding to $\varphi=0.40,0.48,0.51$, and $0.56$, respectively. The blue solid circles indicate the experimental results and the black solid lines are obtained by the numerical solutions of the floating model Eq. (4). The gray area around the solid line indicates the possible solutions within the errors of estimated $\eta$. The black solid lines in large $U_{0}$ are the guide lines $\tilde{F}_{\mathrm{max}}\propto U_{0}^{3/2}$ as expected from Eq. (6).
As explained in Sec. II we can trace the velocity and acceleration of the projectile. By using Eq. (1) we can evaluate the drag force acting on the projectile. Figures 4 depict the time evolutions of the velocity and the drag force acting on the projectile. As shown in Fig. 4 the projectile slows down after the impact by the drag force $F_{D}$ where $F_{D}$ takes the maximum value $F_{\mathrm{max}}$ at $t_{\mathrm{max}}$. In these figures, the dots are obtained in the experiments, and the solid lines are the fitting curves using the solution of Eq.(3). We choose $\eta$ as a fitting parameter to recover experimental $\tilde{F}_{\mathrm{max}}$ in the floating The fitting is performed by obtaining $\eta$ to give the observed in the experiment. Surprisingly, the solution of the floating model Eq. (3) with a fitting parameter $\eta:=3\pi\eta_{{\rm eff}}a_{\rm I}^{3/2}/(g^{1/2}m_{\rm I})$ gives reasonable agreement with the experimental results, at least, for $t<t_{\mathrm{max}}\approx 2.1$ ms corresponding to $F_{\mathrm{max}}$, although there is some discrepancy between the solution of the floating model and the experimental results for $t>t_{\mathrm{max}}$, in which an elastic force might be involved. Here we focus only on the early stage dynamics for $t\leq t_{\mathrm{max}}$ in this paper, the late stage falls outside the scope of the present study. Thus, this result suggests that the floating model is valid to describe the motion of the projectile in the early stage, for $t\leq t_{\mathrm{max}}$, $u_{0}=3.1$ m/s, and $\phi=0.56$.
Figure 5 plots of $\tilde{F}_{\mathrm{max}}$ against $U_{0}$ for various packing fractions, $\phi=0.52,0.53,0.54$, and $0.56$. The blue solid circles indicate experimental estimations of $\tilde{F}_{\mathrm{max}}$, and the black solid lines are the numerical solutions of Eq. (4), where the dimensionless effective viscosity $\eta$ is treated as a fitting parameter with its error bar. The gray areas around the solid line indicate the possible solution within the error of estimated $\eta$. It is readily observed that the theoretical curvesclosely overlap with the experimental data points including their error bars. This agreement is particularly pronounced at higher packing fractions ($\phi=0.54$ and $0.56$), although the theoretical predictions successfully capture the overall trend of the experimental data even at lower packing fractions ($\phi=0.52$ and $0.53$). We also confirm the asymptotic behavior predicted from the floating model as $\tilde{F}_{\mathrm{max}}\sim U_{0}^{3/2}$ for $U_{0}\gg 1$ in Eq. 6 [11, 12]. The solid lines in Fig. 5 indicate the power exponents of the asymptotic solution, which agrees with the slope from the experimental data points. These results indicate that the experimental data across the entire range from $\phi=0.52$ to $\phi=0.56$ are well described by the solution of the floating model given in Eq. (4). Unfortunately, we cannot get definite results for $t_{\mathrm{max}}$, where $t_{\mathrm{max}}$ weakly depends on $U_{0}$. This tendency was reported even in the simulation [11]. We need further investigation on $t_{\mathrm{max}}$.
Figure 6: Plot of the effective viscosity $\eta:=3\pi\eta_{{\rm eff}}a_{\rm I}^{3/2}/(g^{1/2}m_{\rm I})$ against $\phi$. The red vertical lines express error bars of the estimated $\eta$.
Figure 6 exhibits how the dimensionless effective viscosity $\eta$ depends on $\phi$. The figure indicates that $\eta$ weakly depends on $\phi$ in the range $\phi=0.52-0.56$, where $\eta$ slightly increases with $\phi$. This result is unexpected because the viscosity in usual densely packed suspensions increases rapidly as the density approaches the jamming point or the random closed packing value. We will discuss this insensitivity of $\eta$ to $\phi$ in Sec. V.
V Discussion
As presented in the previous section, we found that the floating model provides a good description of our experimental results. In particular, the behavior of the maximum drag force, $\tilde{F}_{\mathrm{max}}$, as a function of the impact speed $U_{0}$ shows reasonable agreement with the theoretical curves. These results validate several simplifications and assumptions made in the floating model, namely: the approximation of buoyancy forces (see Appendix B), the neglect of elastic forces, and the use of Stokes drag when focusing on the early stage of the dynamics. It is important to note that the elastic force becomes significant during the late stage of the dynamics, although its contribution remains much smaller than that of the viscous drag force during the early stage of impact [12].
Surprisingly, the assumption of Stokes drag remains valid, even though it is typically justified only for suspensions of small particles with sufficiently low Reynolds numbers. Nevertheless, this assumption is self-consistent, as explained below. We introduce an effective Reynolds number defined as $Re_{\mathrm{eff}}:=\rho_{\mathrm{f}}U_{0}(d_{\mathrm{max}}+d_{\mathrm{min}})/(2\eta_{\mathrm{eff}})$. Using our experimental parameters, $Re_{\mathrm{eff}}$ ranges from 0.47 to 4.44 at $\phi=0.54$, and from 0.58 to 5.37 at $\phi=0.56$, which are small enough to justify the use of the Stokes approximation. These results support the use of the Stokes drag in the floating model. The apparent viscosity $\eta_{\mathrm{a}}:=\eta_{\mathrm{eff}}/\eta_{0}$, where $\eta_{0}$ is the viscosity of water, ranges from 5675 to 6853. These large values also support the Stokes approximation for the mean-field flow of the suspension liquid.
To evaluate potential sidewall effects due to the quasi-2D container used in the experiment, we also examine a cubic container (Appendix A). The experimental results for $F_{\mathrm{max}}$ versus $U_{0}$ in the cubic container (see Fig.7) are almost indistinguishable from those in Fig.5. This confirms the robustness of the floating model, regardless of the container shape.
Although several models have been proposed to describe impact processes in suspensions [4, 8, 9, 10, 11, 12], most of these models include an elastic term to capture the rebound behavior observed in the late stage of the impact. In contrast, the floating model is particularly well-suited for describing the early-stage dynamics, where viscous effects dominate. Therefore, we conclude that the floating model accurately captures the early-stage dynamics in dense suspensions of millimeter-sized particles.
As shown in Fig. 6, the effective viscosity $\eta$ exhibits only weak dependence on the packing fraction $\phi$. This insensitivity can be attributed to two main factors. First, the actual packing fractions are higher than the nominal ones due to imperfect density matching; for instance, $\varphi=0.40$, 0.48, 0.51, and 0.56 correspond to $\phi=0.52$, 0.53, 0.54, and 0.56, respectively. As a result, our experiments cover only a narrow range of $\phi$. Second, empirical relations suggest that the increase in effective viscosity is not substantial within this narrow range, especially under the assumption that the viscosity diverges at the random close packing fraction $\phi_{\mathrm{rcp}}\approx 0.64$. For example, if we apply the sedimentation theory used for $\phi\leq 0.50$ [23], the ratio of effective viscosity between $\phi=0.56$ and $\phi=0.52$ is about 1.3. Even using empirical relations for denser, sheared non-Brownian suspensions [24], the ratio would be about 2.0. These estimates suggest that the increase in effective viscosity over our observed range is not significant, assuming divergence at $\phi_{\mathrm{rcp}}$, though a detailed quantitative justification remains an open question.
VI Conclusions
In summary, this study demonstrates the applicability of the floating model [9, 11, 12] to suspensions containing millimeter-sized particles, particularly in capturing the relationship between the maximum drag force, $F_{\mathrm{max}}$, and the impact speed, $u_{0}$. Our experimental results confirm the scaling law $F_{\mathrm{max}}\sim u_{0}^{3/2}$ for $u_{0}\gg\sqrt{ga_{\rm I}}$, in agreement with theoretical predictions.
These findings support the validity of previous lattice Boltzmann method (LBM) simulations [11, 12] conducted on sparse suspensions, despite notable differences in the size ratio between the projectile and the suspended particles. Surprisingly, our results suggest that Stokes flow may effectively describe fluid-particle interactions even for large particles, under certain conditions.
Looking ahead, a deeper understanding of why the floating model applies to suspensions of millimeter-sized particles remains an important direction for future research. Although this study primarily focuses on early-stage dynamics, future work should explore the role of elastic forces in the late stages of impact, so-called the IIH as indicated by previous experimental [8] and theoretical [11, 12] studies.
Acknowledgment
We thank the useful comments by Frédéric van Wijland, and Pradipto. HM thanks Mahesh Bandi, and members of the Non-equilibrium Physics Unit at OIST, and Yutaro Motoori for their useful advice and comments. This work is partially supported by Grants-in-Aid of MEXT, Japan for Scientific Research, Grant Nos. JP21H01006 and JP24K17022, and the Kyoto University Foundation. ChatGPT was used for writing assistance and program debugging.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A The dynamical impact of experiments on the cubic cell
In this appendix, we examine a cube container ($W$, $D$, and $H$ are, respectively, given by 94 mm, 94 mm, and 90 mm) to evaluate the possible influence of the side wall effects from the quasi-2D container used in the main text (see Fig. 7). We use the packing fraction $\phi=0.56$ using $N=1197$. The resulting relationship of $F_{\mathrm{max}}$ versus $U_{0}$ behavior closely matches that observed in Fig. 5. Thus, we conclude the robustness of the floating model regardless of container shape or boundary effects.
Figure 7: The plots between the dimensionless maximum force $\tilde{F}_{\mathrm{max}}:=F_{\mathrm{max}}/m_{\rm I}g$ and impact-speed $U_{0}:=u_{0}/\sqrt{a_{\rm I}g}$ for the cubic cell in the packing fraction $\phi=0.56$. The red solid circles indicate the experimental results and the black solid lines are obtained by the numerical solutions of the floating model Eq. (4), where $\eta=12.0\pm 1.2$. The gray area around the solid line indicates the possible solutions within the errors of the estimated $\eta$. The black solid lines in large $U_{0}$ are the guide lines $\tilde{F}_{\mathrm{max}}\propto U_{0}^{3/2}$ as expected from Eq. (6).
Appendix B The expression of the buoyancy force
If a projectile with the density $\rho_{\rm I}$ is located in a fluid with density $\rho_{\rm f}(\neq\rho_{\rm I})$, the motion of the projectile is affected by the buoyancy force. The expression of the buoyancy force must depend on the immersed volume of the projectile in the fluid, where the immersed volume $V(z)$ of the projectile at the location $z$ (see Fig. 2 in the main text) is expressed as
[…]
$=$ $\frac{\pi a_{\rm I}^{3}}{3}\left[3\left(\frac{z}{a_{\rm I}}\right)^{2}-\left(\frac{|z|}{a_{\rm I}}\right)^{3}\right].$
This expression is valid for $-2a_{\rm I}\leq z\leq 0$.
Figure 8: Comparison between the numerical solutions of the exact expression of the buoyancy force Eq. (8) and the approximate expression Eq. (1), where we plot the time evolution of position a) and their difference of two expressions b).
Thus, the buoyancy force is given by
$F_{\rm B}$ $=$ $-g\{\rho_{\rm I}V_{\mathrm{I}}-\rho_{\rm f}V\left(z\right)\}$ (8)
$=$ $-m_{\rm I}g\left\{1-\frac{\rho_{\rm f}}{4\ \rho_{\rm I}}\left[3\left(\frac{z}{a_{\rm I}}\right)^{2}-\left(\frac{|z|}{a_{\rm I}}\right)^{3}\right]\right\}$
[…]
Appendix C Exact solution of floating model
As shown in Refs. [11, 12], the floating model Eq. (4) with Eq. (1) can be solved exactly. The explicit expression of $Z(\tau)$ is given by
$Z\left(\tau\right)=-\left[\frac{4\left(1-\xi\right)}{\eta^{2}}\right]^{1/3}\frac{-{\rm Ai}^{\prime}\left(\Xi\right){\rm Bi}^{\prime}\left(\Lambda\right)+{\rm Ai}^{\prime}\left(\Lambda\right){\rm Bi}^{\prime}\left(\Xi\right)}{{\rm Ai}^{\prime}\left(\Lambda\right){\rm Bi}\left(\Xi\right)+{\rm Ai}\left(\Xi\right){\rm Bi}^{\prime}\left(\Lambda\right)},$ (9)
[…]